OpenStudy (anonymous):

The half-life of iodine-131 is 8.10 days. How long will it take for three-fourths of a sample of iodine-131 to decay? can someone explain to me how to do half-life problems, please?

OpenStudy (aaronq):

For a first order reaction, you can use: \(\Large t_{1/2}=\dfrac{ln2}{k}\), to find k, the decay constant. After you can use the exponential growth/decay equation: \(\Large A_{t}=A_0*e^{-kt}\), where \(A_{t}\) is the amount after t time elapsed, \(A_0\) is the initial amount. t is time, and k is the decay constant (you found above).

OpenStudy (anonymous):

what is In2? im stilll confused?

OpenStudy (anonymous):

can you show me how the problem is done?

OpenStudy (aaronq):

"ln" is the natural logarithm

OpenStudy (anonymous):

oh wowzers

OpenStudy (aaronq):

lol okay, so plug in your half life into the first equation, it should lok ike this: 8.10 = \(\dfrac{ln2}{k}\) solve for k

OpenStudy (aaronq):

use a calculator, obvsiously.

OpenStudy (anonymous):

2.44?

OpenStudy (aaronq):

ln2=0.693 0.693/8.1=0.008557=k

OpenStudy (anonymous):

is that just common knowledge that ln = 0.693? lol

OpenStudy (aaronq):

nope, you have to use your calculator ln(2) lol Theres an easier method for this question, but it won't work for every question of this type.

OpenStudy (anonymous):

ohh well then dont tell me the easier way

OpenStudy (anonymous):

is the answer what k equals?

OpenStudy (aaronq):

no, thats just the decay constant for the next equation. you would plug in your values into this: \(\large A_{t}=A_0*e^{-kt}\) where \(A_{t}\) is the amount after t time elapsed, \(A_0\) is the initial amount. t is time, and k is the decay constant (you found above). can you try to figure out which values go where?

OpenStudy (anonymous):

would At be 8.10 days and Ao would be 3/4?

OpenStudy (aaronq):

no, 8.10 days is the half-life \((t_{1/2})\). we used it in the first part and we don't need to use it again. Ao is the initial amount, so 1 At is the amount you have after 3/4 of the amount decayed, so what is At?

OpenStudy (anonymous):

im just going to take a stab in the dark and say 10.9? i did 181 divided by 3/4

OpenStudy (aaronq):

where did you get 181 from :P lol

OpenStudy (anonymous):

iodine -131? haha

OpenStudy (anonymous):

im still confused as to what A equals

OpenStudy (aaronq):

131 is just the mass of the isotope, thats not really relevant to the question here. so Ao is the initial amount, that we set as 1 arbitrarily (we could've said it was any number, but we choose 1 to make it easy). so At is the amount after "t" time elapsed, they said 3/4 of the sample decayed, so At=1-3/4=1/4 cool?

OpenStudy (anonymous):

got it!

OpenStudy (anonymous):

so 1/4=1 times e ^ -0.0008557 ?

OpenStudy (aaronq):

okay, so we plug everything in: \(\large \dfrac{1}{4}=(1)*e^{-0.008557\color{red}t}\) now, we only need to solve for t. do you know how to solve logarithms?

OpenStudy (anonymous):

ahh no :( sorry.. dont i just plug it in my calulator?

OpenStudy (aaronq):

yeah, but you first need to rearrange it using logarithm rules. Unless you have a really good calculator.

OpenStudy (anonymous):

how do i arrange it?

OpenStudy (aaronq):

by the way i wrote too many zeroes for k, k=0.08557 \(\dfrac{1}{4}=(1)*e^{-0.08557t}\) take "ln" of both sides \(ln(\dfrac{1}{4})=ln(e^{-0.08557t})\) using logarithm rules the exponent can come down \(ln(\dfrac{1}{4})=-0.08557t*ln(e)\) \(ln(e)=1\; because \;ln=log_e, \; so \;ln_e(e)=1\) were left with: \(ln(\dfrac{1}{4})=-0.08557t\) from here just isolate t \(\dfrac{ln(\dfrac{1}{4})}{-0.08557}=t\) plug into your calculator

OpenStudy (anonymous):

16.20?

OpenStudy (aaronq):

yep. thats right. if you notice the easier way would've been finding how many half-lives it takes to get to 1/4 of the original. if you have 1 to start with after one half life you have 1/2 then after other half life you have 1/4 2 * half-life=2*8.1 days = 16.2 days

OpenStudy (anonymous):

hahah oh wow so much easier that way! but now i know for other problems(: thanks so much i appreciate your time!(:

OpenStudy (aaronq):

haha yeah, it's a lot easier but it won't work if they give you random numbers like, find how long it takes to have 0.18491923 of the original. With the method i showed you you CAN solve that. lol no problem !